Find the percentage of increase or decrease for the following:
2 to 15
20 to 13
Thank you very much!

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

15/2=7.5 OR 750% INCREASE
(20-13)/20 OR 7/20=.35 OR 35% DECREASE

Answer by stanbon(75887)

Find the percentage of increase or decrease for the following:
2 to 15 is an increase of 15/2=7.5 =750/100= 750%
20 to 13 is a decrease of 1-13/20= 100%-65%=35%
Cheers,
Stan H.